OpenStudy (anonymous):

When simplifying (3-7/x)/(x+5/6x), do you times all the bits of the fraction by 6x to get rid of the fractions in the fraction, or do you times the top by x and the bottom by 6x?

OpenStudy (lalaly):

is it\[\huge{\frac{\frac{3-7}{x}}{\frac{x+5}{6x}}}\]

OpenStudy (anonymous):

No, its (3-(7/x)) / (x + (5/6x))

OpenStudy (lalaly):

\[\huge{\frac{3-\frac{7}{x}}{x+\frac{5}{6x}}}\]

OpenStudy (lalaly):

?

OpenStudy (anonymous):

yeah, sorry. Thanks for clarifying.

OpenStudy (lalaly):

:D its ok you multiply top and bottom by 6x

OpenStudy (anonymous):

won't that leave the top with a fraction? Wouldn't the top be 18x- (42x/6) or something?

OpenStudy (lalaly):

it will be like\[\frac{18x-42}{6x^2+30}\] then y can divide them all by 6 u get \[\frac{6x-7}{x^2+5}\]

OpenStudy (lalaly):

sorry the first one is 3x not 6x

OpenStudy (anonymous):

I see! Thank you!

OpenStudy (lalaly):

ur welcome :D

OpenStudy (lalaly):

i made a miistake :S

OpenStudy (lalaly):

its supposed to be\[\frac{18x-42}{6x^2+5}\]because the 6x cancels

OpenStudy (anonymous):

Ohh, right. Okay. You're right. I missed that too. :) Thanks.

OpenStudy (lalaly):

im sorry:)

OpenStudy (anonymous):

That's okay! I'm grateful for the help!

OpenStudy (lalaly):

=),=)

OpenStudy (anonymous):

So is it -4/x+5, I wonder?